OpenStudy (anonymous):

Simplify. Express the product as a radical expression. (x^3)^(3/4)

OpenStudy (anonymous):

x^9/4

OpenStudy (anonymous):

\[\sqrt[4]{x^{9}}\]

hero (hero):

Both answers above are correct, but no one stated any general formulas nor provided any explanations. I'm greatly disappointed.

OpenStudy (anonymous):

Can't that be simplified to \[x ^{2}\sqrt[4]{x}\]

hero (hero):

I was just about to mention that.

OpenStudy (anonymous):

\[(x^{m})^{n} = x^{m \times n }\] m = 3, n = 3/4 m*n = 3*(3/4) = 9/4\[(x^{m})^{n} = x^{m \times n }\]

OpenStudy (anonymous):

Haha, ok xD Thanks guys!